→ 2.0^{5000.0} x^{y} √

Command  Key Binding
Move forward past an expression ESC-right or ESC-Ctl-F
Move backward past an expression ESC-left or ESC-Ctl-B
Select forward expression  ESC-shift-right or ESC-Ctl-Shift-F
Select backward expression  ESC-shift-left or ESC-Ctl-Shift-B
Move outside an expression  ESC-up or ESC-Ctl-U
Move inside an expression  ESC-down or ESC-Ctl-D